OpenStudy (anonymous):

The midpoint of a line segment with end points as (-1, 5) and (x2, 7) is (-4, 6). What is the value of x2? -14 -2 -7 -3

OpenStudy (anonymous):

Use the midpoint formula

OpenStudy (anonymous):

Thus, -1 + x2 = 2(-4) -1 + x2 = -8 x2 = -8 + 1 x2 = -7

OpenStudy (anonymous):

Thank you!